# GURecon: Learning Detailed 3D Geometric Uncertainties for Neural Surface Reconstruction

### v2

@@ Abstract
Neural surface representation has demonstrated remarkable success in the areas of novel view synthesis and 3D reconstruction. However, assessing the geometric quality of 3D reconstructions in the absence of ground truth mesh remains a significant challenge, due to its rendering-based optimization process and entangled learning of appearance and geometry with photometric losses. In this paper, we present a novel framework, i.e, GURecon, which establishes a geometric uncertainty field for the neural surface based on geometric consistency. Different from existing methods that rely on rendering-based measurement, GURecon models a continuous 3D uncertainty field for the reconstructed surface, and is learned by an online distillation approach without introducing real geometric information for supervision. Moreover, in order to mitigate the interference of illumination on geometric consistency, a decoupled field is learned and exploited to finetune the uncertainty field. Experiments on various datasets demonstrate the superiority of GURecon in modeling 3D geometric uncertainty, as well as its plug-and-play extension to various neural surface representations and improvement on downstream tasks such as incremental reconstruction. The code and supplementary material are available on the project website: https://zju3dv.github.io/GURecon/.
@@ 1 Introduction
Image-based 3D reconstruction is a long-standing problem in computer vision with a wide range of applications like AR/VR, autonomous driving, digital heritage preservation, etc. Recently, learning-based methods have attracted much attention with the development of neural radiance representations like Neural Radiance Fields (NeRF) (Mildenhall et al. 2021). Unlike traditional methods, NeRF and its variants (e.g, NSVF (Liu et al. 2020), NeuS (Wang et al. 2021)) encode scene geometry and appearance with neural networks, which can be optimized by leveraging the differentiable rendering given a set of calibrated images.
@@
In this paper, we present a novel framework, i.e, GURecon, which is able to learn detailed 3D geometric uncertainty for neural surface reconstruction as shown in Fig. 1. Different from existing methods relying on rendering-based pixel-wise uncertainty measurement, GURecon directly models the 3D uncertainty for surface points and ensures consistency over viewpoints. However, designing such a system is nontrivial. Without ground truth geometric supervision (e.g, input depth), it is difficult to model the geometric uncertainty just based on photometric error between the rendered and input images. This is because, as with previous methods, the neural radiance field tends to overfit the input images in sparse viewpoint settings, resulting in minor photometric errors but significant geometric errors, just as the ambiguity problem highlighted in NeRF++ (Zhang et al. 2020).
Motivated by the traditional multi-view stereo (Hu and Mordohai 2012; Schönberger et al. 2016) where photometric consistency is widely used to assess the confidence of reconstructed geometry, we employ the multi-view consistency as a cue to quantify the quality of reconstruction. We compute the patch-based warping consistency of surface points projected onto the input images, and utilize it as a pseudo label of geometric accuracy to supervise a continuous geometric uncertainty field based on a novel online distillation approach. We consider the estimated uncertainty derived from such pseudo label as epistemic uncertainty, which reflects the geometric confidence of the model per-scene trained with given images (i.e, reconstruction error), and serves as a reference identifying areas where reconstruction is inadequate and unreliable.
@@
  • We present a novel framework, i.e, GURecon, to quantify geometric uncertainty for neural surface reconstruction.
@@ 3 Method
In this paper, we introduce a novel framework, i.e, GURecon, which enables accurate geometric uncertainty estimation for various neural surface representations without GT geometric information for supervision. As shown in Fig. 2, with given posed images, we learn a render field and an SDF field through differentiable rendering. As the training progresses, with the currently learned geometry field, we first utilize a root-finding method to identify the zero-crossing points intersected with the implicit surface and calculate the multi-view consistency of these points as pseudo supervision to guide the learning of geometric uncertainty (Sec. 3.2). Then we present a novel online distillation method that simultaneously learns a spatially continuous uncertainty field with other fields in a self-supervised manner by utilizing the multi-view consistency as pseudo ground truth labels (Sec. 3.3). In order to overcome the interference caused by view-dependent factors in the calculation of multi-view consistency, we propose to simultaneously learn additional decoupled fields and exploit them to fine-tune the geometric uncertainty field (Sec. 3.4).
@@ 3.2 Patch-based Multi-view Consistency
where $\mathbf{n}^{\prime}$ is the normal computed from the gradient of the SDF values, i.e, $\nabla f_{s}(\mathbf{p}^{\prime})$, and $d$ is the distance to the origin of the coordinate system. Then, the homography warping matrix H can be constructed based on the local plane and enables the mutual projection of image patches between viewpoints as the following:
@@ 4.1 Uncertainty Quantification
Metrics. In the experiment for uncertainty quantification, following previous methods (Shen et al. 2022; Goli et al. 2023), we calculate the Area Under Sparsification Error (AUSE), a widely used metric to assess the quality of model uncertainty. Given the predicted depth error and predicted uncertainty of each pixel in the test image, we gradually remove the top $t\%(t=1\sim 100)$ pixels according to two criteria: once based on GT depth error, once based on predicted uncertainty, and compute the average depth error for the remaining pixels. The area between the curves obtained from the two criteria is the AUSE, which reflects the correlation between predicted uncertainty and actual depth error. In addition to calculating AUSE based on the Mean Absolute depth Error ($\Delta$MAE) and Mean Squared Error ($\Delta$MSE), we also compute the AUSE based on the closest distance of each point to the ground truth geometry as the 3D geometric error, i.e, $\mathrm{AUSE}_{\mathrm{3D}}$, which reflects geometric uncertainty from a 3D perspective. Finally, we evaluate the accuracy of surfaces reconstructed by different methods using the Chamfer Distance (CD). Since the scales among scenes are inconsistent in both BlendedMVS and TNT datasets, we uniformly normalize the scenes to fit within the unit sphere to compute geometric metrics.
@@ 5 Conclusion
In this paper, we introduce GURecon, a novel approach for learning a 3D geometric uncertainty field for neural surface models. Unlike existing methods that model rendering-based pixel-wise uncertainty, the proposed GURecon exploits the multi-view consistency to accurately model the geometric uncertainty. Moreover, additional decoupled fields are learned for robust uncertainty estimation. Comprehensive experiments have demonstrated our superior performance compared to existing methods. As Fig. 10 shows, while our approach works well in small-scale textureless regions, its performance is limited in extreme scenarios with large textureless areas (e.g, white walls), where high-level semantic information can be incorporated in future work.
6 Acknowledgments
We express our gratitude to all the anonymous reviewers for their professional and constructive comments. This work was partially supported by the NSFC (No. 62102356), Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.

### v3

@@ Abstract
Neural surface representation has demonstrated remarkable success in the areas of novel view synthesis and 3D reconstruction. However, assessing the geometric quality of 3D reconstructions in the absence of ground truth mesh remains a significant challenge, due to its rendering-based optimization process and entangled learning of appearance and geometry with photometric losses. In this paper, we present a novel framework, GURecon, which establishes a geometric uncertainty field for the neural surface based on geometric consistency. Different from existing methods that rely on rendering-based measurement, GURecon models a continuous 3D uncertainty field for the reconstructed surface, and is learned by an online distillation approach without introducing real geometric information for supervision. Moreover, in order to mitigate the interference of illumination on geometric consistency, a decoupled field is learned and exploited to finetune the uncertainty field. Experiments on various datasets demonstrate the superiority of GURecon in modeling 3D geometric uncertainty, as well as its plug-and-play extension to various neural surface representations and improvement on downstream tasks such as incremental reconstruction.
@@ 1 Introduction
Image-based 3D reconstruction is a long-standing problem in computer vision with a wide range of applications like AR/VR, autonomous driving, digital heritage preservation, etc. Recently, learning-based methods have attracted much attention with the development of neural radiance representations like Neural Radiance Fields (NeRF) (Mildenhall et al. 2021). Unlike traditional methods, NeRF and its variants (e.g., NSVF (Liu et al. 2020), NeuS (Wang et al. 2021)) encode scene geometry and appearance with neural networks, which can be optimized by leveraging the differentiable rendering given a set of calibrated images.
@@
In this paper, we present a novel framework, i.e., GURecon, which is able to learn detailed 3D geometric uncertainty for neural surface reconstruction as shown in Fig. 1. Different from existing methods relying on rendering-based pixel-wise uncertainty measurement, GURecon directly models the 3D uncertainty for surface points and ensures consistency over viewpoints. However, designing such a system is nontrivial. Without ground truth geometric supervision (e.g., input depth), it is difficult to model the geometric uncertainty just based on photometric error between the rendered and input images. This is because, as with previous methods, the neural radiance field tends to overfit the input images in sparse viewpoint settings, resulting in minor photometric errors but significant geometric errors, just as the ambiguity problem highlighted in NeRF++ (Zhang et al. 2020).
Motivated by the traditional multi-view stereo (Hu and Mordohai 2012; Schönberger et al. 2016) where photometric consistency is widely used to assess the confidence of reconstructed geometry, we employ the multi-view consistency as a cue to quantify the quality of reconstruction. We compute the patch-based warping consistency of surface points projected onto the input images, and utilize it as a pseudo label of geometric accuracy to supervise a continuous geometric uncertainty field based on a novel online distillation approach. We consider the estimated uncertainty derived from such pseudo label as epistemic uncertainty, which reflects the geometric confidence of the model per-scene trained with given images (i.e., reconstruction error), and serves as a reference identifying areas where reconstruction is inadequate and unreliable.
@@
  • We present a novel framework, i.e., GURecon, to quantify geometric uncertainty for neural surface reconstruction.
@@ 3 Method
In this paper, we introduce a novel framework, i.e., GURecon, which enables accurate geometric uncertainty estimation for various neural surface representations without GT geometric information for supervision. As shown in Fig. 2, with given posed images, we learn a render field and an SDF field through differentiable rendering. As the training progresses, with the currently learned geometry field, we first utilize a root-finding method to identify the zero-crossing points intersected with the implicit surface and calculate the multi-view consistency of these points as pseudo supervision to guide the learning of geometric uncertainty (Sec. 3.2). Then we present a novel online distillation method that simultaneously learns a spatially continuous uncertainty field with other fields in a self-supervised manner by utilizing the multi-view consistency as pseudo ground truth labels (Sec. 3.3). In order to overcome the interference caused by view-dependent factors in the calculation of multi-view consistency, we propose to simultaneously learn additional decoupled fields and exploit them to fine-tune the geometric uncertainty field (Sec. 3.4).
@@ 3.2 Patch-based Multi-view Consistency
Surface Interaction Retrieval. The primary step is to identify the surface points of the neural representation. Following the existing works (Fu et al. 2022; Oechsle, Peng, and Geiger 2021), root finding is a widely used method to locate the intersection with the neural surface. As our approach is based on SDF representation, and the SDF values of sampling points along the ray are precomputed for volume rendering, we employ linear interpolation to locate the zero-crossing points $\mathcal{T}$ as follows:
@@
where $s_{i}$ is the SDF value of $\mathbf{p}(t_{i})$, i.e., $f_{s}(\mathbf{p}(t_{i}))$, and $\hat{t}_{i}$ is the ray segment of the zero-crossing point.
@@
where $\mathbf{n}^{\prime}$ is the normal computed from the gradient of the SDF values, i.e., $\nabla f_{s}(\mathbf{p}^{\prime})$, and $d$ is the distance to the origin of the coordinate system. Then, the homography warping matrix H can be constructed based on the local plane and enables the mutual projection of image patches between viewpoints as the following:
@@ 4.1 Uncertainty Quantification
Metrics. In the experiment for uncertainty quantification, following previous methods (Shen et al. 2022; Goli et al. 2023), we calculate the Area Under Sparsification Error (AUSE), a widely used metric to assess the quality of model uncertainty. Given the predicted depth error and predicted uncertainty of each pixel in the test image, we gradually remove the top $t\%(t=1\sim 100)$ pixels according to two criteria: once based on GT depth error, once based on predicted uncertainty, and compute the average depth error for the remaining pixels. The area between the curves obtained from the two criteria is the AUSE, which reflects the correlation between predicted uncertainty and actual depth error. In addition to calculating AUSE based on the Mean Absolute depth Error ($\Delta$MAE) and Mean Squared Error ($\Delta$MSE), we also compute the AUSE based on the closest distance of each point to the ground truth geometry as the 3D geometric error, i.e., $\mathrm{AUSE}_{\mathrm{3D}}$, which reflects geometric uncertainty from a 3D perspective. Finally, we evaluate the accuracy of surfaces reconstructed by different methods using the Chamfer Distance (CD). Since the scales among scenes are inconsistent in both BlendedMVS and TNT datasets, we uniformly normalize the scenes to fit within the unit sphere to compute geometric metrics.
@@ 5 Conclusion
In this paper, we introduce GURecon, a novel approach for learning a 3D geometric uncertainty field for neural surface models. Unlike existing methods that model rendering-based pixel-wise uncertainty, the proposed GURecon exploits the multi-view consistency to accurately model the geometric uncertainty. Moreover, additional decoupled fields are learned for robust uncertainty estimation. Comprehensive experiments have demonstrated our superior performance compared to existing methods. While our approach works well in small-scale textureless regions, its performance is limited in extreme scenarios with large textureless areas (e.g., white walls), where high-level semantic information can be incorporated in future work.
Acknowledgments
We express our gratitude to all the anonymous reviewers for their professional and constructive comments. This work was partially supported by the NSFC (No. 62441222), Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.